OpenStudy (anonymous):

Calculate the speed of a proton that is accelerated from rest through a potential difference of 198 V. Answer in units of m/s

OpenStudy (anonymous):

energy = (1/2) m v^2 = 198 electron-Volts = 198 eV 1 eV = 1.6x10^-19 J proton mass = m = 1.67 x10^27 kg from these you can calculate V^2 and thus v, m/s

OpenStudy (whpalmer4):

That should be \(m = 1.672622*10^{\color{red}{-}27} \text { kg}\)

OpenStudy (anonymous):

Thank you. Especially important is that - sign, in 10^-27 kg. Sorry.

OpenStudy (anonymous):

Thank you :)